OpenStudy (anonymous):

Identify the 35th term of an arithmetic sequence where a1 = -7 and a18 = 95.

OpenStudy (anonymous):

Ive been working pretty diligently on this problem, and yet cant seem to get the right answer.

OpenStudy (anonymous):

a18=a1+(n-1)*d 95=-7+17d 17d=102 d=6 a35=a1+(n-1)*d=-7+34*6=197

OpenStudy (anonymous):

wha? omg it that easy! where in gods name have i pulled out my other answers from? that....jesus....haha i feel like such a dummy

OpenStudy (anonymous):

Well, I have seen this before, but I don't remember how to solve it... Sorry...

OpenStudy (anonymous):

lol indigoeve

OpenStudy (anonymous):

I did it just that way....and somehow came up with a totally different answer than 6. I got like...14.6.....HOW?

OpenStudy (anonymous):

gosh thank you so much guys haha just got me past a brain fart

OpenStudy (anonymous):

Well, that's how you know the man that sold you your calculator can't be trusted... damn that man

OpenStudy (anonymous):

your so very welcome hun ask me any question and i could proably help lol

OpenStudy (anonymous):

b-but. I spent a lot of money on my calculator. (Admittedly I was just using my computer calculator and was just too lazy to get up and go get my graphing calc lol)

OpenStudy (anonymous):

lol lazinees get u now were but look at justin bieber lol

OpenStudy (anonymous):

Hey by the way, if you ever need help with anything just tag me. I help mostly with biology. And history, English, sometimes math, etc. etc.

OpenStudy (anonymous):

i mostly math and chemistry

OpenStudy (anonymous):

Do either of you do chem? I'm bomb at lit and history but chem and math....eh not so much. Numbers arent really my forte

OpenStudy (anonymous):

I can also help with physics and chemistry. :3

OpenStudy (anonymous):

OH YOU DO :D sweet beans guys! You both are my new best friends.

OpenStudy (anonymous):

Yay! Friends! I wonder what it feels like...

OpenStudy (anonymous):

yay bf lol nice ur welcome sweety

OpenStudy (anonymous):

gosh i need all the help i can get. You'll have to deal with the occasional stupidity but other than that, It'll be great! xD

OpenStudy (anonymous):

im use to it i tried helping this blonde thats the only time i quit

OpenStudy (anonymous):

well lucky for you I'mm a natural brown haha

OpenStudy (anonymous):

hah i can tell im looking at ur pic wich i like alot

OpenStudy (anonymous):

Would you guys be able to check my work on another problem of this sort to make sure I got it?

OpenStudy (anonymous):

thank you. My hair is actually red at the moment

OpenStudy (anonymous):

Sure. Math?

OpenStudy (anonymous):

sure hun

OpenStudy (anonymous):

yes sir

OpenStudy (anonymous):

It'll be up in a sec

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

i have medals flying at me help lol

OpenStudy (anonymous):

Hahaha! Medal spam! Success! X3